Solve this Let a = i + 2 j + 3k and b = 3 i + j find the unit vector in the direction of
a+b?
 

Grade:12

1 Answers

KUNCHAM SAMPATH
8 years ago
Given a = i + 2 j + 3k and b = 3i + j
a + b = i + 2 j + 3k + 3i + j
= 4 i + 3 j + 3k
∴ The unit vector in the direction of
=± 4 i + 3 j + 3k
   16 + 9 + 9
=± 
a+b
|a+b|
4 i + 3 j + 3k
| 4 i + 3 j + 3k |
=± 
a+b=±
4 i + 3 j + 3k
34